OpenStudy (anonymous):

Can someone please help me more then what I have got... Solve by the elimination method 0.3x-0.2y=4 0.4y+0.5y=1 What is the order pair?

OpenStudy (anonymous):

The equations right? Would think the bottom one would need an x somewhere...

OpenStudy (anonymous):

Well you need to either get "x" or "y" alike so they can cancel

OpenStudy (anonymous):

0.4 is x sorry

OpenStudy (anonymous):

Ok so do what ebezzy said. Try multiplying the equations by a constant i.e a number so you can get either x or y to cancel when you add them.

OpenStudy (anonymous):

0.4(0.3x-0.2y)=4 = (0.12x-0.8y)= 1.6 -0.3(0.4y+0.5y)=1= (-0.12-0.15y)=-0.30 -.95y=1.3 so y = -1.3 now solve for x

OpenStudy (anonymous):

I am very sorry my pc is messing up really bad

OpenStudy (anonymous):

I am getting a decimal with over 10 numbers please help more